# Images are Achilles’ Heel of Alignment: Exploiting Visual Vulnerabilities for Jailbreaking Multimodal Large Language Models

### v2

@@ 1 Introduction
Figure 1: An example to show the influence of the visual modality on harmlessness alignment of Gemini Pro Vision. The harmful information is highlighted in red.
However, it still lacks of deep understanding of how safety concerns occur in MLLMs and how they might be different from those in LLMs. Considering this issue, this work aims to systematically analyze the sourcing factors that violate the harmlessness alignment of MLLMs. We conduct detailed empirical studies on representative MLLMs, specifically investigating their performance on harmful instructions accompanied by images. Our findings are threefold: (1) Images can be backdoors for the harmlessness alignment of MLLMs. The inclusion of images in the input can significantly increase the harmfulness ratio of MLLMs’ outputs; (2) Cross-modal finetuning undermines the alignment abilities of the backbone LLM for a given MLLM. The more parameters that are finetuned, the more severe the disruption is; (3) The harmfulness of MLLMs’ responses is positively correlated with the harmfulness of the image content. These findings reveal that the visual modality introduces additional alignment vulnerabilities in MLLMs, which can be exploited to further jailbreak these models.
@@ 2.1 Evaluation Data Collection
To evaluate the harmlessness alignment of MLLMs, we collect a dataset comprising 750 harmful instructions across 5 scenarios. Each instruction includes a harmful keyword or key phrase and is paired with a harmful image related to the keyword or key phrase. We present the collection process below and show the pipeline in appendix.
First, based on existing harmful scenarios of LLMs [14], we select five representative ones that are related to the visual information in the real world: (1) Violence, Aiding and Abetting, Incitement; (2) Financial Crime, Property Crime, Theft; (3) Privacy Violation; (4) Self-Harm; and (5) Animal Abuse. For simplicity, these categories are referred to as Violence, Financial, Privacy, Self-Harm, and Animal, respectively. Next, we adopt GPT-4 to generate 50 keywords for each of the above harmful categories, and then synthesize three harmful but distinct instructions based on each keyword. The prompts employed for the above process are documented in appendix. In this way, we guarantee that each instruction includes only one harmful element (i.e., the keyword or the key phrase), which could be accurately depicted by an image. Thus, we can pair each instruction with a corresponding real-world image that is relevant to the harmful keyword/phrase. Specifically, we first retrieve five images from Google using each keyword/phrase as the query, and then employ CLIP ViT-L/14 [25] to select the image that best matches the semantic representation of the keyword/phrase.
@@ 2.2 Evaluation Settings
Table 1: The evaluation results of representative MLLMs on the dataset we collected. (Train) represents the cross-modal finetuning strategies of MLLMs. Animal, Financial, Privacy, Self-Harm, and Violence represent the ASR of MLLMs on instructions from these categories. Average represents the average ASR across all categories. $+$ and -- represents the change of ASR compared to the Backbone setting.
@@ 2.3 Evaluation Results
Images can be alignment backdoors of MLLMs. When comparing the performance of each model under the Backbone and Text-only settings, the harmlessness alignment of MLLMs does not significantly deviate from that of their backbone LLMs, and even exhibit enhanced defense capability, e.g., LLaVA-1.5 (--2.80%) and Gemini ProV (--9.65%). However, once adding images, regardless of whether their contents are harmful or not, the ASR results of MLLMs would be greatly improved even under the Blank setting using harmless images, e.g., LLaVA-1.5 ($+$25.20%) and MiniGPT-v2 ($+$33.87%). It indicates that images can be the alignment backdoor of MLLMs, which would undermine MLLMs’ capability of defending against harmful text input.
More parameters tuned, less alignment left. By examining the performance of open-sourced MLLMs, we notice that their alignment performance is highly relevant to their training strategies during cross-modal finetuning. Generally, for the same backbone model, the more parameters are optimized during finetuning, the harmlessness alignment would be affected more significantly. For example, under the Toxic setting, MiniGPT-4, with the fixed backbone LLM, is more robust than the LoRA-finetuned MiniGPT-v2, achieving lower ASR results (23.47% vs. 43.87%). Similarly, the full-parameter finetuned LLaVA-1.5 also generates more harmful responses than LoRA-finetuned LLaVA-1.5L (68.13% vs. 62.80%). The reason may be that the cross-modal finetuning process would hurt the harmlessness alignment of the backbone LLMs.
@@ 3.2 Amplifying Image Harmfulness with LLMs
The whole procedure of image harmfulness optimization is presented in appendix. We leverage the harmfulness of the caption as the proxy for the image’s harmfulness, as the harmfulness of text is easier to quantify than images. We consider an iterative process to generate harmful images. We first ask ChatGPT to modify the original instruction $t$ into an initial image generation prompt $p_{0}$ and generate an initial image. At step $k$, the caption model $\mathcal{C}$ generates a caption $c_{k}$ for the target image $i_{\text{opt}}^{k}$. Subsequently, the judging model $\mathcal{J}$ assesses the image’s harmfulness with a score, $s_{k}$, on a scale from 1 to 10, where a higher score indicates greater harmfulness. $\mathcal{J}$ also explains the reason for its score in $exp_{k}$. All this information ($p_{k}$, $c_{k}$, $s_{k}$, and $exp_{k}$) is appended to the conversation history $h$ and sent to the attacker model $\mathcal{A}$. which first suggests improvements to the prompt and then generates the refined image generation prompt $p_{k+1}$. The refined prompt is then used to generate a new image $i_{\text{opt}}^{k+1}$. The above process circulates until reaching the maximum iteration turn $K$, which is set to 5. In practice, we choose GPT-4-0613 as both the attacker and judging model by utilizing different system prompts (as presented in the appendix). For captioning and image generation, we adopt LLaVA-1.5 and PixArt-$\alpha$ [7], respectively.
@@ 4.3 Further Analyses
From the charts, we notice that most jailbroken cases are due to the conflicts between the general instruction following abilities of MLLMs and their harmlessness alignment, while captioning and OCR abilities also occupy considerable proportions. The results suggest that the cross-modal finetuning may impose a kind of “inverse alignment tax” on MLLMs, which improves their multimodal abilities while impairing the harmlessness alignment. Therefore, to enhance the harmlessness alignment of MLLMs, future work could consider adding more adversarial samples that consist of harmful instructions and images during the cross-modal finetuning process, which ensures MLLMs defend attacks from the image side while maintaining multimodal abilities.
@@ 5 Related Work
Alignment refers to the process of fine-tuning LLMs with annotations based on human preferences, to ensure that the generated responses of LLMs are Helpful, Honest, and Harmless, a.k.a., the 3H principle [2]. The harmlessness alignment of LLMs has attracted extensive research attention. To evaluate the robustness of LLMs to harmful instructions, researchers employ red teaming to benchmark the safety degree of MLLMs [13, 11]. Additionally, some studies further explore the harmlessness alignment of LLMs with jailbreaking methods, either in white-box or black-box attacks. In white-box attacks, attackers have full access to the target LLMs, allowing them to calculate and backpropagate gradients from customized harmful outputs to inputs, which guides the target LLMs to generate harmful content [33, 41, 29]. In contrast, in black-box attacks, attackers have no access to the model’s parameters or gradient information, and only launch attacks through the prompting interface [35] or LLM attacker [5, 34]. Our work mainly focuses on extending the jailbreaking research from LLMs to MLLMs, which aims to enhance the robustness and alignment of MLLMs.
Multimodal LLMs.
By leveraging the excellent capacities of LLMs, the community has witnessed the emergence of several studies on MLLMs. Specifically, LLaVA [18] and MiniGPT-4 [40], which integrate an LLM, a vision encoder, and a cross-modal connector, introduce two-stage training strategies, namely cross-modal representation aligning and visual instruction tuning. Following this procedure, many open-source MLLMs have been released, by improving the multimodal performance with high-resolution and stronger LLMs [17], more pre-training corpus [3], and specially pre-training tasks [3, 6].
@@
Similar to LLMs, MLLMs also suffer from malicious prompts and likely generate harmful content [24, 38]. To explore the harmfulness of MLLMs, some studies try to prob the potential harmfulness of MLLMs with benchmarks, while others employ white- and black-box attacks on MLLMs with adversarial images and prompts. However, existing benchmarks mainly focus on the effectiveness capacities of MLLMs, lacking sufficient consideration of harmlessness evaluation [10, 15, 20]. To bridge this gap, recent studies have developed a red teaming attack benchmark covering different scenarios to systematically analyze the harmlessness of MLLMs [32, 19, 16]. Similar to LLMs, existing work on jailbreaking MLLMs can be categorized into white- and black-box attacks. For white-box attacks on MLLMs, existing work [24, 26, 4, 27, 9, 30] can be further classified into two categories according to attacking objectives: input images and visual embeddings. For input images, recent studies generate adversarial images with constraints of a harmful response set [24, 26, 9] or utilizing a teacher-forcing optimization approach [4]. For visual embeddings, Shayegani et al. [27] generates adversarial images that look harmless but are similar to the embeddings of harmful images, thereby bypassing harmful content filters. In contrast, recent work in black-box attacks jailbreak the harmlessness alignment of MLLMs by employing techniques such as system prompt attacks [36], transferring harmful information into text-oriented images [12], generating adversarial images with surrogate models [39], and maximum likelihood-based jailbreaking [21]. In our work, we first investigate whether the visual input influences the harmlessness alignment of MLLMs and then incorporate promoting engineering, white- and black-box attacks on MLLMs.
@@ 6 Conclusion
In this paper, we conducted a comprehensive empirical analysis of the harmlessness alignment of MLLMs, specifically examining the visual vulnerabilities for jailbreak. Our findings revealed that images pose significant vulnerabilities in the alignment of MLLMs: the presence of images, the cross-modal finetuning process, and the harmfulness of images all contribute to an increased propensity for MLLMs to generate harmful responses. Furthermore, we introduced HADES, a novel jailbreaking approach that hides and amplifies the harmfulness of textual instructions using meticulously crafted images. Extensive experiments have demonstrated that HADES is capable of effectively jailbreaking both open- and closed-source MLLMs. In summary, our work has presented strong evidence that the visual modality poses the alignment vulnerability of MLLMs, underscoring the urgent need for further exploration into cross-modal alignment.
For future work, we will consider improving our jailbreak strategies to more effectively jailbreak well-aligned MLLMs like GPT-4V. And we will devise cross-modal alignment methods to enhance the harmlessness alignment of MLLMs.
@@ 0.A.1 Experimental Setting
To avoid data leakage issues, we only utilize the first 50% data from HADES to generate instructions and utilize the remaining data for evaluation. We totally collect 2286 instructions. We shuffle all these instructions while ensuring that each negative instruction is followed by its positive counterpart. This method aims to teach MLLMs to differentiate between harmful and harmless instructions by contrasting them, thereby learning which instructions should be followed. Subsequently, we adopt these instructions to finetune LLaVA-1.5 using LoRA. The resulting LoRA is named as contrastive harmlessness LoRA.
@@ 0.A.2 Results and Analysis
Appendix 0.B Data Collection Pipeline
@@
Appendix 0.C Additional Jailbreak Cases
@@
Appendix 0.D ChatGPT Prompt for Harmful Instruction Generation
@@
Appendix 0.E Prompts for the Attacker and Judging Model
In this section, we present the system prompt for attacker and judging model in Fig. 11 and Fig. 12, respectively.
@@ Appendix 0.E Prompts for the Attacker and Judging Model
Appendix 0.F Pseudo Code for Image Harmfulness Optimization

### v3

@@ 1 Introduction
Figure 1: An example to show the influence of the visual modality on harmlessness alignment of Gemini Pro Vision. The harmful information is highlighted in red
However, it still lacks a deep understanding of how safety concerns occur in MLLMs and how they might be different from those in LLMs. Considering this issue, this work aims to systematically analyze the sourcing factors that violate the harmlessness alignment of MLLMs. We conduct detailed empirical studies on representative MLLMs, specifically investigating their performance on harmful instructions accompanied by images. Our findings are threefold: (1) Images can be backdoors for the harmlessness alignment of MLLMs. The inclusion of images in the input can significantly increase the harmfulness ratio of MLLMs’ outputs; (2) Cross-modal fine-tuning undermines the alignment abilities of the backbone LLM for a given MLLM. The more parameters that are fine-tuned, the more severe the disruption is; (3) The harmfulness of MLLMs’ responses is positively correlated with the harmfulness of the image content. These findings reveal that the visual modality introduces additional alignment vulnerabilities in MLLMs, which can be exploited to further jailbreak these models.
@@ 2.1 Evaluation Data Collection
To evaluate the harmlessness alignment of MLLMs, we collect a dataset comprising 750 harmful instructions across 5 scenarios. Each instruction includes a harmful keyword or key phrase and is paired with a harmful image related to the keyword or key phrase. We present the collection process below and show the pipeline in the supplementary materials.
First, based on existing harmful scenarios of LLMs [12], we select five representative ones that are related to the visual information in the real world: (1) Violence, Aiding and Abetting, Incitement; (2) Financial Crime, Property Crime, Theft; (3) Privacy Violation; (4) Self-Harm; and (5) Animal Abuse. For simplicity, these categories are referred to as Violence, Financial, Privacy, Self-Harm, and Animal, respectively. Next, we adopt GPT-4 to generate 50 keywords for each of the above harmful categories, and then synthesize three harmful but distinct instructions based on each keyword. The prompts employed for the above process are documented in the supplementary materials. In this way, we guarantee that each instruction includes only one harmful element (i.e., the keyword or the key phrase), which could be accurately depicted by an image. Thus, we can pair each instruction with a corresponding real-world image that is relevant to the harmful keyword/phrase. Specifically, we first retrieve five images from Google using each keyword/phrase as the query, and then employ CLIP ViT-L/14 [21] to select the image that best matches the semantic representation of the keyword/phrase.
@@ 2.2 Evaluation Settings
Table 1: The evaluation results of representative MLLMs on the dataset we collected. (Train) represents the cross-modal fine-tuning strategies of MLLMs. Animal, Financial, Privacy, Self-Harm, and Violence represent the ASR of MLLMs on instructions from these categories. Average represents the average ASR across all categories. $+$ and -- represents the change of ASR compared to the Backbone setting.
@@ 2.3 Evaluation Results
Images can be alignment backdoors of MLLMs. When comparing the performance of each model under the Backbone and Text-only settings, the harmlessness alignment of MLLMs does not significantly deviate from that of their backbone LLMs, and even exhibit enhanced defense capability, e.g., LLaVA-1.5 (--2.80%) and Gemini ProV (--7.50%). However, once adding images, regardless of whether their contents are harmful or not, the ASR results of MLLMs would be greatly improved even under the Blank setting using harmless images, e.g., LLaVA-1.5 ($+$25.20%) and MiniGPT-v2 ($+$33.87%). It indicates that images can be the alignment backdoor of MLLMs, which would undermine MLLMs’ capability of defending against harmful text input.
More parameters tuned, less alignment left. By examining the performance of open-sourced MLLMs, we notice that their alignment performance is highly relevant to their training strategies during cross-modal fine-tuning. Generally, for the same backbone model, the more parameters are optimized during fine-tuning, the harmlessness alignment would be affected more significantly. For example, under the Toxic setting, MiniGPT-4, with the fixed backbone LLM, is more robust than the LoRA-fine-tuned MiniGPT-v2, achieving lower ASR results (23.47% vs. 43.87%). Similarly, the full-parameter fine-tuned LLaVA-1.5 also generates more harmful responses than LoRA-fine-tuned LLaVA-1.5L (68.13% vs. 62.80%). The reason may be that the cross-modal fine-tuning process would hurt the harmlessness alignment of the backbone LLMs.
@@ 3.2 Amplifying Image Harmfulness with LLMs
The whole procedure of image harmfulness optimization is presented in the supplementary materials. We leverage the harmfulness of the caption as the proxy for the image’s harmfulness, as the harmfulness of text is easier to quantify than images. We consider an iterative process to generate harmful images. We first ask ChatGPT to modify the original instruction $t$ into an initial image generation prompt $p_{0}$ and generate an initial image. At step $k$, the caption model $\mathcal{C}$ generates a caption $c_{k}$ for the target image $i_{\text{opt}}^{k}$. Subsequently, the judging model $\mathcal{J}$ assesses the image’s harmfulness with a score, $s_{k}$, on a scale from 1 to 10, where a higher score indicates greater harmfulness. $\mathcal{J}$ also explains the reason for its score in $exp_{k}$. All this information ($p_{k}$, $c_{k}$, $s_{k}$, and $exp_{k}$) is appended to the conversation history $h$ and sent to the attacker model $\mathcal{A}$. which first suggests improvements to the prompt and then generates the refined image generation prompt $p_{k+1}$. The refined prompt is then used to generate a new image $i_{\text{opt}}^{k+1}$. The above process circulates until reaching the maximum iteration turn $K$, which is set to 5. In practice, we choose GPT-4-0613 as both the attacker and judging model by utilizing different system prompts (as presented in the supplementary materials). For captioning and image generation, we adopt LLaVA-1.5 and PixArt-$\alpha$ [6], respectively.
@@ 4.3 Further Analyses
From the charts, we notice that most jailbroken cases are due to the conflicts between the general instruction following abilities of MLLMs and their harmlessness alignment, while captioning and OCR abilities also occupy considerable proportions. The results suggest that the cross-modal fine-tuning may impose a kind of “inverse alignment tax” on MLLMs, which improves their multimodal abilities while impairing the harmlessness alignment. Therefore, to enhance the harmlessness alignment of MLLMs, future work could consider adding more adversarial samples that consist of harmful instructions and images during the cross-modal fine-tuning process, which ensures MLLMs defend attacks from the image side while maintaining multimodal abilities.
@@ 5 Related Work
Alignment involves fine-tuning LLMs with human-preferred annotations to ensure responses are Helpful, Honest, and Harmless (the 3H principle [2]). Among them, the harmlessness alignment of LLMs has attracted extensive research attention. To evaluate the robustness of LLMs to harmful instructions, researchers employ red teaming to benchmark the safety degree of LLMs [11, 9]. Additionally, some studies further explore the harmlessness alignment of LLMs with jailbreaking methods. Some adopt white-box attacks, which utilize the model gradients to customize adversarial inputs [28, 36, 25], while black-box attacks typically launch attacks through manually or automatically devised prompts [30, 4, 29]. Our work mainly focuses on extending the jailbreaking research from LLMs to MLLMs, which aims to enhance the robustness and alignment of MLLMs.
@@
By utilizing LLMs as backbones, MLLMs also inherit their alignment vulnerabilities. To explore the harmlessness alignment of MLLMs, several benchmarks are proposed to prob the potential harmfulness of MLLMs under different scenarios [27, 16, 13]. Some other work employ different jailbreak methods to further evaluate the adversarial robustness of MLLMs using white- or black-box methods. White-box methods mainly attack the input images or their embeddings of MLLMs. For input images, recent studies generate adversarial images with constraints of a harmful response set [20, 22, 8] or utilizing a teacher-forcing optimization approach [3]. For visual embeddings, Shayegani et al. [23] generates adversarial images that look harmless but are similar to the embeddings of harmful images, thereby bypassing harmful content filters. In contrast, recent work in black-box attacks jailbreak MLLMs by employing techniques such as system prompt attacks [31], transferring harmful information into text-oriented images [10], generating adversarial images with surrogate models [34], and maximum likelihood-based jailbreaking [17]. In our work, we first investigate how the visual input influences the harmlessness alignment of MLLMs, then propose a jailbreak methods incorporate both white- and black-box methods.
@@ 6 Conclusion
In this paper, we conducted a comprehensive empirical analysis of the harmlessness alignment of MLLMs, specifically examining the visual vulnerabilities for jailbreak. Our findings revealed that images pose significant vulnerabilities in the alignment of MLLMs: the presence of images, the cross-modal fine-tuning process, and the harmfulness of images all contribute to an increased propensity for MLLMs to generate harmful responses. Furthermore, we introduced HADES, a novel jailbreaking approach that hides and amplifies the harmfulness of textual instructions using meticulously crafted images. Extensive experiments have demonstrated that HADES can effectively jailbreak both open- and closed-source MLLMs. In summary, our work has presented strong evidence that the visual modality poses the alignment vulnerability of MLLMs, underscoring the urgent need for further exploration into cross-modal alignment. In future work, we will develop cross-modal training strategies to improve the harmlessness alignment of MLLMs.
6.0.1 Societal Impacts.
Our work aims to highlight the alignment vulnerabilities of existing MLLMs. We hope our jailbreak attempt can guide subsequent researchers in developing safer MLLMs. However, we acknowledge that certain elements of our research, such as harmful instructions and images, may have negative societal impacts. To minimize these negative effects, we have implemented several measures, including adding warnings in the abstract and placing safety statements on the dataset homepage. Furthermore, in the supplementary materials, we preliminarily explore how to use the HADES data to fine-tune MLLMs to enhance their safety alignment. Overall, we believe with these efforts, the positive contributions of our work outweigh its potential negative impacts.
Acknowledgement
This work was partially supported by National Natural Science Foundation of China under Grant No. 62222215, Beijing Natural Science Foundation under Grant No. L233008 and 4222027. Xin Zhao is the corresponding author.
@@ 0.A.1 Experimental Setting
To avoid data leakage issues, we only utilize the first 50% data from HADES to generate instructions and utilize the remaining data for evaluation. We totally collect 2286 instructions. We shuffle all these instructions while ensuring that each negative instruction is followed by its positive counterpart. This method aims to teach MLLMs to differentiate between harmful and harmless instructions by contrasting them, thereby learning which instructions should be followed. Subsequently, we adopt these instructions to fine-tune LLaVA-1.5 using LoRA. The resulting LoRA is named as contrastive harmlessness LoRA.
@@ 0.A.2 Results and Analysis
Appendix 0.B Comparison with other jailbreak methods.
We compared HADES with two other representative jailbreak methods for MLLMs, represented as Adversarial [22] and Compositional [23], respectively. We implemented these methods against LLaVA-1.5 on our collected dataset. The results are presented in Tab. 4, where HADES achieves the highest ASR across all categories.
Table 4: ASR of new baselines and HADES on LLaVA-1.5.
Methods Animal Financial Privacy Self-Harm Violence
Adversarial [22] 74.67 84.00 89.33 80.67 86.67
Compositional [23] 54.67 78.00 81.33 48.00 84.00
HADES 83.33 89.33 94.67 89.33 94.67
Appendix 0.C Comparison between Beaver-dam-7B and human annotation.
To ensure the reliability of adopting LLMs to evaluate the harmfulness of models’ responses, we engaged three human annotators to annotate the output of LLaVA-1.5, according to the original annotation documents from BeaverTails. As the results in Tab. 5, Beaver-dam-7B is highly consistent with humans in assessing the harmfulness of MLLMs’ responses.
Table 5: The evaluation results of Beaver-dam-7B and human annotators.
Annotator Animal Financial Privacy Self-Harm Violence
Beaver-dam-7B 83.33 89.33 94.67 89.33 94.67
Human 84.00 88.67 91.33 88.00 92.67
Appendix 0.D Data Collection Pipeline
@@
Appendix 0.E Additional Jailbreak Cases
@@
Appendix 0.F ChatGPT Prompt for Harmful Instruction Generation
@@
Appendix 0.G Prompts for the Attacker and Judging Model
In this section, we present the system prompt for the attacker and judging model in Fig. 11 and Fig. 12, respectively.
@@ Appendix 0.G Prompts for the Attacker and Judging Model
Appendix 0.H Pseudo Code for Image Harmfulness Optimization